OpenStudy (anonymous):

Find trignometric function of 5 pi /6

ganeshie8 (ganeshie8):

u may use below : \(\sin (\pi-x) = \sin (x)\) \(\cos (\pi-x) = -\cos (x)\)

OpenStudy (anonymous):

i got it i made a unit circle

ganeshie8 (ganeshie8):

thats clever ^

OpenStudy (anonymous):

\[\sin \frac{ 5π }{ 6 }=\sin \frac{ (6-1)π }{ 6 }=\sin \left( π-\frac{ π }{ 6 } \right)=\sin \left( \frac{ π }{ 6 } \right)=1/2\] \[\cos ^{2}x+\sin^{2}x=1 <=> cosx=\pm \sqrt{1-\sin^{2}x} =\pm \sqrt{1-\frac{ 1 }{ 4 }}=\pm \frac{ \sqrt{3} }{ 2 }\] \[tanx=\frac{ sinx }{ cosx }= \frac{ \frac{ 1 }{ 2 } }{ \frac{\pm \sqrt{3} }{ 2 } }=\pm \frac{ \sqrt{3} }{ 3 }\] \[cotx=\tan^{-1} x=\pm \sqrt{3}\] ... .. .

ganeshie8 (ganeshie8):

Nice :)